The difference of twice one number and five time another number is 11. The first number is three time the second number increased by 6. What are the numbers?
Found 2 solutions by macston, fractalier:
Answer by macston(5194)

F=first number; S=second number
.
F=3S-6
.
2F-5S=11
2(3S-6)-5S=11
6S-12-5S=11
S=23
.
ANSWER 1: The second number is 23.
.
F=3S-6=3(23)-6=69-6=63
ANSWER 2: The first number is 63.
.
CHECK:
2F-5S=11
2(63)-5(23)=11
126-115=11
11=11

Answer by fractalier(6550)

Call the first and second numbers, x and y. Then we write
2x - 5y = 11 and
x = 3y + 6
Now substitute the second equation into the first and get
2(3y+6) - 5y = 11
6y + 12 - 5y = 11
y + 12 = 11
y = -1 and
x = 3(-1) + 6 = 3
Your numbers are 3 and -1.